# Effect of recommending users and opinions on the network connectivity and idea generation process

### v1

@@ 1   Introduction
With the increasing popularity of social networks, they have proven to be a great tool to communicate opinions, spread the information and influence the opinions [1, 2]. However, the spread of information and opinions over social networks is not unimpeded and free. It is controlled and moderated for several reasons and by several parties including but not limited to social media algorithms, political and business entities [3, 4, 5]. We have seen a massive development in the field of recommendation systems over the last decades. Recommendation systems have emerged as essential tools for numerous industries, providing personalized and relevant recommendations to individual users, which is critical for generating revenue and gaining a competitive edge in e-commerce platforms [6, 7]. Recommendation systems find application in a diverse range of fields, including e-commerce, entertainment, research, personalized health and fitness, and more [6, 8, 9].
@@
These studies contribute significantly to our understanding of the complex relationship between recommender systems and opinion dynamics in social networks. We contribute to this area with extending one of recent works by Sayama [38] by incorporating the concept of recommender systems in it. In his work, Sayama investigates the role of conformity, homophily and attention to novel opinions on the network dynamics utilizing a numerical simulation model. We modify existing model to include several recommendation and filtering strategies, to generate and collect opinions and to measure the opinion eccentricity [39] to better understand the effect of recommendation systems on network dynamics and idea generation process. By including recommender systems in the simulation model, we can examine how these systems impact the network’s polarization, fragmentation, and the diversity of ideas. Our research extends the current understanding of opinion dynamics by incorporating recommender systems, which have become an integral part of social networks. We aim to identify potential risks associated with recommendation and filtering strategies and inform the design of more responsible and effective algorithms.
@@ 2.2   Updated model
where $x_{max}=1$, $x_{min}=10^{-6}$ and $\alpha=3$ and $u$ is a uniform random number.
@@
At each time step both node states and edge weights undergo changes as well. However, to minimize computational complexity, we update the idea state and node weight only after every 100 time steps. Node states and edge weights co-evolve over time according to a modified version of the co-evolution mechanism used in the original model. The co-evolution mechanism is driven by homophily, attention to novelty, conformity, and random fluctuation, as in the original model. However, in our model, exposure to new opinions or users is controlled by the recommendation strategy being used.
@@
Here $\overline{X^{exposed}_{i}}$ is the weighted average of exposed opinions to node $i$ where the weight is of edge from author of opinion to user $i$; $\overline{X^{recent}_{j}}$ is the average of recent 10 opinions generated by node $j$; $c$, $h$, and $a$ are the parameters determining strength of social conformity, homophily, and attention to novelty, respectively. $\mathcal{E}$ represents a random fluctuation. $F_{h}$ and $F_{a}$ are functions determining the edge weight change as follows:
@@ 2.3   Recommendation strategies
  • 20 Nearest opinions (NO): In this recommendation strategy, each user is exposed to 20 opinions from the pool of 100 new opinions generated in the whole network that are most similar to their idea state. Each user also updates the weight of their edge with the authors of these 20 opinions.
@@ 3   Experiments
We ran two sets of simulation runs to investigate the role of recommendation systems. The parameter combination used for both sets of experiments are as follows:
@@ 3.1   Network dynamics
The first set of experiments we conduct is to examine the relationship between recommendation strategy and network evolution. In the first set, we ran 15 simulation runs for each combination of the above parameters. In the first set, we ran 15 simulation runs for each combination of the parameters. The outcome measures we calculated in these experiments are maximum modularity of network and maximum standard deviation of average community state. We run each simulation until there are 3,000 opinions in the network and after every 100 new opinions we calculate the modularity and standard deviation of average community idea states. The Louvain’s method [40] is used to identify network fragmentation and communities in the network.
@@ 3.2   Idea generation process
In the second set of experiments, we investigate the impact of recommendation strategies on eccentric idea generation. For each parameter combination, we calculate eccentricity of each opinion generated in the network and the overall distribution of eccentricities is observed. Eccentricity is the measure of strangeness of an opinion relative to the social neighborhood [39]. Eccentricity has been proven an important metric in idea generation process to understand the social impact [41]. Since eccentricity calculation is computationally complex, we only ran 3 simulation runs for each combination of parameters. Eccentricity calculation is performed as follows. The norm or center of a user’s knowledgebase is calculated by taking the weighted average of neighbors’ recent opinions. Distance of a new opinion from the center is the eccentricity of the new opinion. We use $L^{2}$ norm to find the distance between opinion and the norm.

### v2

@@ 1   Introduction
With the increasing popularity of social networks, they have proven to be a great tool to communicate opinions, spread the information and influence opinions [1, 2]. However, the spread of information and opinions over social networks is not unimpeded and free. It is controlled and moderated for several reasons and by several parties including but not limited to social media algorithms, political and business entities [3, 4, 5]. We have seen a massive development in the field of recommendation systems over the last decades. Recommendation systems have emerged as essential tools for numerous industries, providing personalized and relevant recommendations to individual users, which is critical for generating revenue and gaining a competitive edge in e-commerce platforms [6, 7]. Recommendation systems find application in a diverse range of fields, including e-commerce, entertainment, research, personalized health and fitness, and more [6, 8, 9].
@@
These studies contribute significantly to our understanding of the complex relationship between recommender systems and opinion dynamics in social networks. We extend one of the recent works of Sayama [38] by incorporating the concept of recommender systems in it. In his work, Sayama investigates the role of conformity, homophily and attention to novel opinions on the network dynamics utilizing a numerical simulation model. We modify the existing model to include several recommendation and filtering strategies, to generate and collect opinions and to measure the opinion eccentricity [39] to better understand the effect of recommendation systems on network dynamics and idea generation process. By including recommender systems in the simulation model, we can examine how these systems impact the network’s polarization, fragmentation, and the diversity of ideas. Our research extends the current understanding of opinion dynamics by incorporating recommender systems, which have become an integral part of social networks. We aim to identify potential risks associated with recommendation and filtering strategies and inform the design of more responsible and effective algorithms.
@@ 2.2   Updated model
where $x_{max}=1$, $x_{min}=10^{-6}$, $\alpha=3$, and $u$ is a uniform random number. $\alpha$ value was chosen as $3$ since it represents one end of the spectrum within the typical range of $\alpha$ values in social systems ( $2\leq\alpha\leq 3$). The goal of this study is to analyze effects of recommendation systems; uniform and power law distributions are selected to exhibit the robustness of our findings. To precisely model any particular social network, the value of $\alpha$ should be adjusted accordingly.
@@
At each time step both node states and edge weights undergo changes as well. However, to minimize computational complexity, we update the idea state and node weight only after every 100 time steps. Node states and edge weights co-evolve over time according to a modified version of the co-evolution mechanism used in the original model. The co-evolution mechanism is driven by homophily, attention to novelty, conformity, and random fluctuation, as in the original model. However, in our model, exposure to new opinions or users is controlled by the recommendation strategy being used. $\overline{X^{exposed}_{i}}$ is the weighted average of all the opinions node $i$ is exposed to( $\overline{X^{exposed}_{i}}=\frac{\sum_{k}w_{ik}\cdot X_{k}}{\sum_{k}w_{ik}}$, where $X_{k}$ is one of the opinions user $i$ is exposed to, $k$ is its author, and $w_{ik}$ is the weight of the edge from user $k$ to user $i$ ). $\overline{X^{recent}_{j}}$ is the average of recent 10 opinions generated by node $j$ ( $\overline{X^{recent}_{j}}=\frac{\sum_{m}X_{m}}{N}$, where $X_{m}$ represents one of the recent opinions generated by node $j$, $N$ is the number of recent opinions considered, which is 10 in our study).
The co-evolution takes place as follows:
@@
Here $\overline{X^{exposed}_{i}}$ is the weighted average of all the opinions node $i$ is exposed to. Weighted average is calculated using the weight of edge from author of an opinion to user $i$; ; $c$, $h$, and $a$ are the parameters determining strength of social conformity, homophily, and attention to novelty, respectively. $\mathcal{E}$ represents a random fluctuation. $F_{h}$ and $F_{a}$ are functions determining the edge weight change as follows:
@@ 2.3   Recommendation strategies
  • 20 Nearest opinions (NO): In this recommendation strategy, each user is exposed to 20 opinions from the pool of 100 new opinions generated in the whole network that are most similar to their idea state. Each user also updates the weight of their edge with the authors of these 20 opinions. Nearness is calculated using the euclidean distance or $L_{2}$ norm between opinion vectors or idea states according to the strategy.
@@ 3   Experiments
We ran two sets of simulations to investigate the role of recommendation systems. The parameter combination used for both sets of experiments are as follows:
@@ 3.1   Network dynamics
The first set of experiments we conduct is to examine the relationship between recommendation strategy and network evolution. In the first set, we conducted 15 simulation runs for each combination of the above parameters. The outcome measures we calculated in these experiments are maximum modularity of network and maximum standard deviation of average community state. We run each simulation until there are 3,000 opinions in the network and after every 100 new opinions we calculate the modularity and standard deviation of average community idea states. The Louvain’s method [40] is used to identify network fragmentation and communities in the network.
@@ 3.2   Idea generation process
In the second set of experiments, we investigate the impact of recommendation strategies on eccentric idea generation. For each parameter combination, we calculate eccentricity of each opinion generated in the network and the overall distribution of eccentricities is observed. Eccentricity is the measure of strangeness of an opinion relative to the social neighborhood [39]. Eccentricity has been proven an important metric in idea generation process to understand the social impact [41]. Since eccentricity calculation is computationally complex, we only ran 3 simulation runs for each combination of parameters. Eccentricity calculation is performed as follows. The norm or center of a user’s knowledgebase is calculated by taking the weighted average of neighbors’ recent opinions. Distance of a new opinion from the center is the eccentricity of the new opinion. We use $L^{2}$ norm to find the distance between opinion and the center.
@@ 4.1   Effect of recommendation strategies on network dynamics
(a) Low homophily, high attention to novelty and initial weights follow power law distribution
(b) Low homophily, high attention to novelty and initial weights follow uniform random distribution